OpenStudy (anonymous):

What values for q (0 < q < 2p) satisfy the equation? 2(square root 2) sin q + 2 = 0.

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

^D

OpenStudy (anonymous):

\[2\sqrt{2}\sin(q)+2=0\] \[2\sqrt{2}\sin(q)=-2\] \[\sin(q) = -\frac{2}{2\sqrt{2}}\] \[\sin(q) = -\frac{\sqrt{2}}{2}\] Can you go on from here? :)

OpenStudy (anonymous):

I really don't understand :( I am sorry

OpenStudy (anonymous):

Basically, for what values of q is sin(q) equal to (1/2)*sqrt(2)?